OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. P1 M P2 (12, 4) (2, 6)

OpenStudy (anonymous):

i need to find p2

OpenStudy (anonymous):

12 to 2 is ten left. another ten left is -8 4 to 6 is up 2, another 2 up gives 8 so answer is \((-8,8)\)

OpenStudy (anonymous):

ohhhh... thansk!! how about this one... Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. P1 M P2 (-16, 0) (0, -16) would it be (0,16)

OpenStudy (anonymous):

you can check that it is right, since \((\frac{12-8}{2},\frac{8+4}{2})\) is \((2,6)\)

OpenStudy (anonymous):

ohhh thank u very much!

OpenStudy (anonymous):

are you looking for the midpoint or the endpoint in your second question?

OpenStudy (anonymous):

no. p2

OpenStudy (anonymous):

ok so for first coordinate you need to solve \[\frac{-16+x}{2}=0\] or \[-16+x=0\] \[x=16\] so first coordinate is 16

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

second coordinate solve \[\frac{0+y}{2}=-16\] \[\frac{y}{2}=-16\] \[y=-32\] so your point is \((16,-32)\)

OpenStudy (anonymous):

oh thanks! thanks! appreciate it